Solve by substitution,state whether it is independent,inconsistent,or dependent.;
x - y = 3
3x -2y = 3

Answer by stanbon(57347)

1st: x - y = 3
2nd: 3x -2y = 3
-------------
Multiply 1st by 3 to get:
3rd: 3x - 3y = 9
--------------------
Subtract 2nd from 3rd to get:
-y = 6
y = -6
--------------
Substitute into 1st to solve for "x"
x -(-6) = 3
x+6 = 3
x = -3
----------------
Check the answer in 2nd:
3(-3) - 2(-6) = 3
-9 +12 = 3
3 = 3
-----------------
The system is consistent.
========================
Cheers,
Stan H.